# Learning in Markets with Heterogeneous Agents: Dynamics and Survival of Bayesian vs. No-Regret Learners

### v1

@@ 5 The Regret of Bayesian Learners
In this section, we further analyze the relationship between an agent’s regret to its wealth level and long-term survival in market dynamics, and use this analysis to derive the regret level of Bayesian learners. This relationship depends on the competition. In particular, we saw in Theorem 3.1 that the learners who survive (have a positive expected wealth level in the limit) are only those who maintain a constant regret gap relative to the best competitor in the market.
@@
Note that this agent uses information regarding future realizations of random states, which is not available to real agents in a stochastic environment. Next, we derive the regret level of using the correct distribution of states $q$ as the investment strategy.
@@
Proof.
Consider an agent indexed by $q$ using the state distribution $q$ as its investment strategy; i.e, $\alpha^{q}_{t}=q$ for all $t$, and we denote the seqeunce by $q_{1:T}$. Agent $1$ competes with a magic agent. The regret of the agent playing strategy $q$ is denoted $R^{T}(q_{1:T})$, and denote its expectation by $R=\operatorname{\mathbb{E}}[R^{T}(q_{1:T})]$. On one hand, using the fact that $\hat{q}$ has no regret by definition, from Equation 16 we have
$R^{T}(q_{1:T})=\log(r^{\hat{q}q}_{T})-\log(r^{\hat{q}q}_{0}).$
On the other hand, taking an expectation of Equation 6, we have
$\operatorname{\mathbb{E}}\big{[}\log(r^{\hat{q}q}_{T})\big{]}$ $=$ $\operatorname{\mathbb{E}}\big{[}\sum_{s}\sum_{t}\mathbf{1}_{\{s_{t}=s\}}\log(\frac{\alpha^{\hat{q}}_{s}}{q_{s}})\big{]}+\log(r^{\hat{q}q}_{0})$
$=$ $T\operatorname{\mathbb{E}}\big{[}\sum_{s}\frac{n_{s}}{T}\log(\frac{n_{s}/T}{q_{s}})\big{]}+\log(r^{\hat{q}q}_{0})=T\cdot\operatorname{\mathbb{E}}[I_{\hat{q}}(q)]+\log(r^{\hat{q}q}_{0}).$
where $n_{s}$ is the empirical count of state $s$ until time $T$. And thus, we get
$R=T\cdot\operatorname{\mathbb{E}}\big{[}I_{\hat{q}}(q)\big{]}.$
That is, the expected regret equals the relative entropy (a.k.a. KL divergence) between an empirical distribution and the state distribution scaled by $T$. The question of survival of this agent now hinges on how does $I_{\hat{q}}(q)$ depend on time. Here, we use a result from information theory, showing that the expected KL divergence of the empirical distribution decreases as the alphabet size divided by the sample size [^7] We are interested in the expectation. [68] also provide concentration bounds for this result. (see, e.g., Lemma 13.2 in [70]). In our case, $\operatorname{\mathbb{E}}[I_{\hat{q}}(q)]\equiv\operatorname{\mathbb{E}}[D_{KL}(\hat{q}||q)]=\frac{S-1}{2T}+o(\frac{1}{T})$. Thus, the expected regret $R$ is constant, depending only on the number of states. ∎
We can now also state the following result about a learner competing with a magic agent.
@@
An agent indexed by $q$ using the correct distribution $q$ survives when competing against a magic agent. This is since by Theorem 5.2, the strategy $q$ yields constant expected regret $R$, and so the regret difference compared to $\hat{q}$—which by definition yields zero regret—is constant as well. By Theorem 3.1, this implies that $q$ survives.
Our analyses above and in the preceding sections lead to the following characterization of the regret of Bayesian learners and the survival conditions for other learning agents competing with them in the market.
@@
Proof.
By Theorem 5.2, an agent $n$ using $\alpha^{n}_{t}=q$ for all $t$ incurs constant expected regret. By Proposition 4.2, a Bayesian learner whose support includes the distribution $q$ almost surely survives against such a competitor. Next, Theorem 3.1 shows that if some agent survives against a competitor, then the difference between their regret levels must be bounded by a constant at all times. Therefore, the Bayesian competing against the state distribution $q$ almost surely incurs constant regret. Having established the first point, the second follows directly from Theorem 3.1: Since the Bayesian has constant regret, any competitor who survives must also have constant regret. ∎
Before concluding this section, the following lemma provides an expression for the regret of a player as a function of the entropy relative to strategy $q$, using our notation $R$ from Theorem 5.2 for the expected regret of strategy $q$. This will be useful in the following Section on imperfect Bayesians.
@@
The expected regret of a constant strategy sequence ($\alpha^{n}_{1:T}$ such that $\alpha^{n}_{t}=\alpha^{n}$ for all $t$) is
$R^{T}(\alpha^{n}_{1:T})=T\cdot I_{q}(\alpha^{n})+R.$
Proof.
This follows from the definition of regret as a utility difference, as discussed in Section 3. Specifically, the regret of strategy sequence $\alpha^{n}_{1:T}$ is $R^{T}(\alpha^{n}_{1:T})=U(\alpha^{\hat{q}}_{1:T})-U(\alpha^{n}_{1:T})$, also, $R=\operatorname{\mathbb{E}}[U(\alpha^{\hat{q}}_{1:T})-U(q_{1:T})]$, and denote the regret of $\alpha^{n}_{1:T}$ with respect to the benchmark $q_{1:T}$ as $R^{\alpha}_{q}=U(q_{1:T})-U(\alpha^{n}_{1:T})$. We have $\operatorname{\mathbb{E}}[R^{T}(\alpha^{n}_{1:T})]=\operatorname{\mathbb{E}}[R^{\alpha}_{q}]+R=\operatorname{\mathbb{E}}\Big{[}T\sum_{s}\frac{n_{s}}{T}\log(\frac{q_{s}}{\alpha^{n}_{s}})\Big{]}+R=T\cdot I_{q}(\alpha^{n})+R$. ∎
@@ 6 Imperfect Bayesians
In the preceding sections, we saw that a perfect Bayesian is optimal in the sense that it survives almost surely in the market and drives out any player with regret increasing over time. In this section, we explore a scenario where a Bayesian learner suffers small errors, either in its set of prior distributions or in its update rule. We find that, while perfect Bayesians are optimal, Bayesian learning is also very fragile; for example, it is key that the correct probabilities is one of the model they consider. Specifically, we show that an imperfect Bayesian would eventually vanish from the market, either against the state distribution $q$ or in competition with any player with sub-linear regret. This holds even if the Bayesian’s errors are very small and have zero mean.
@@ 6.1 Bayesian Learning with Inaccurate Priors
Note that this contrasts with the case of Theorem 5.4, where a perfect Bayesian dominates the market in competition with any learners with regret increasing in time.
Proof.
Let $n$ be the index of a Bayesian agent with $q$ not within its support, and let $q^{\prime}\neq q$ be the strategy within the support that is closest to $q$ in relative entropy. By Proposition 4.1, the strategy used by the Bayesian agent, $\alpha^{n}_{t}$, converges almost surely to $q^{\prime}$. Hence, intuitively, for sufficiently large $t$, the strategies remain bounded away from $q$, leading to linear regret from that point onward. In more detail, define $I_{0}=\frac{1}{2}I_{q}(q^{\prime})>0$. In every sample path $i$ (i.e., an infinite sequence, indexed by $i$, of state realizations and the ensuing sequence of $\alpha^{n}_{t}$), except on a set of measure zero, there exists a time $T_{i}$ such that for all $t\geq T_{i}$, we have $I_{q}(\alpha^{n}_{t})>I_{0}$. The regret accumulated up to time $T_{i}$ is constant (possibly positive or negative). The expected regret difference relative to strategy $q$ at $t>T_{i}$ is $\operatorname{\mathbb{E}}[U^{t}(q)-U^{t}(\alpha^{n}_{t})]\geq I_{0}$, and so the expected regret relative to $q$ until time $T$ is at least $I_{0}\cdot(T-T_{i})+const$. By Theorem 5.2, the strategy $q$ incurs only constant regret, so the regret of the Bayesian agent converging to $q^{\prime}$ grows linearly in $T$. Now, consider a no-regret learner. Such a learner has sublinear worst-case regret and thus sublinear expected regret. By Theorem 3.1, the imperfect Bayesian agent who has linear regret vanishes. ∎
@@
Our goal now is to analyze the survival time during which the inaccurate Bayesian still maintains a significant market share, and to understand how this survival time is related to the level of error $\epsilon$. This, of course, also depends on the best competitor and how fast they converge.
To estimate this relation and provide an upper bound on the survival time of an inaccurate Bayesian learner, we consider a player using $q^{\prime}$ with $TV(q^{\prime},q)=\epsilon$ (having already converged to this distribution), competing against a second player playing a dynamic strategy with expected regret level increasing as $f(t)$ that is sub-linear with $t$.
First, we would like to express the first player’s error in terms of entropy (i.e., KL divergence with $q$). We can bound the entropy using Pinsker’s inequality [69] for a lower bound and its inverse for an upper bound, were the latter holds for finite distributions with full support, as in our case. We have (where recall that $\Delta=\min_{s}q_{s}$)
@@
To maximize the survival time, we pick the distribution $q^{\prime}$ to be the one that has the smallest regret, which is when $I_{q}(q^{\prime})$ is the smallest possible: $I_{q}(q^{\prime})=2\epsilon^{2}$. So we have [^8] If the inaccurate agent uses the worst $q^{\prime}$ with error $\epsilon$, the regret calcualtion is similar, but with $\epsilon$ rescaled by a factor of $1/\sqrt{\Delta}$, using the inverse of Pinsker’s inequality: $2\epsilon^{2}\tau/\Delta+R$. The smallest-regret $q^{\prime}$ is used for the bound. $\mathbb{E}[R^{T}(q^{\prime}_{1:T})]=2\epsilon^{2}T+R$. By Equation (16), the regret difference between two players is equal to the log of their wealth ratios plus a constant. By comparing regret levels between the agents, we obtain a bound on the typical survival time $\tau$ up to which the inaccurate player holds, in expectation, more than half the market value; beyond this point, their expected share is less and continues decaying to zero: [^9] An alternative question one may want to ask is, “How accurate should my prior be if I want to hold a share of the market for $T$ steps against the competitor?” For this, one can invert the equation to get $\epsilon=\sqrt{(f(T)-R)/T}$.
@@
Remark: Note that $f(t)$ is the competitor’s actual regret. Regret bounds for known learning algorithms (e.g., bandit algorithms such as UCB [75] or gradient-descent and second-order methods like those described in [56]) are typically worst-case bounds. Estimating the expected regret (or the most likely one) in game dynamics in general, or specifically in our investment scenario, is an interesting open problem.
For the special cases that the competitor has constant regret (e.g., an accurate Bayesian), or the competitor has $\log T$ or $\sqrt{T}$ regret we get the following results.
@@
For the the competitor with logarithmic regret, we get an equation of the form $\tau=c_{1}\cdot\exp(c_{2}\epsilon^{2}\tau)$. For small errors $\epsilon$, by expanding the exponent, we have the following bound.
@@
Finally, in competition against a learner with regret $f(T)=\sqrt{T}$ the survival time is longer.
@@
The interpretation of the above analysis depends, of course, on the time scale we are interested in. When $T$ is large, which is our main focus (e.g., when trade occurs at high frequency or if investments are long-term), only the long-term survival matters. In this case, a Bayesian with an inaccurate prior will eventually vanish in competition with any learner who converges to the truth (such as a regret minimizer). However, when considering transient states as well, it is possible that a Bayesian with a prior that is inaccurate but relatively close to the true distribution could hold a significant market share for some time before eventually vanishing. See Section 7 for an example of such a scenario.
@@ 6.2 Bayesian Learning with Noisy Updates
Next, we consider a different type of imperfect Bayesian learner that does have $q$ in its prior, but in every step performs slightly inaccurate “trembling hand” updates. Also here, we demonstrate that the optimality of Bayesian learning is very fragile, even when the correct distribution lies in the support. To model this, we define a noisy Bayesian learner as one who at each step, either slightly over-weights the current observation or slightly over-weights its current prior, such that, in expectation, both the data and the prior receive the correct weights in every step (i.e., the errors in weight have zero mean).
For concreteness, consider the following scenario of a learner attempting to learn a distribution of states. Suppose there are two states, $s_{t}=0$ with a fixed probability $q\in(0,1)$, and $s_{t}=1$ otherwise. The learner considers two models: $\theta_{a}=q$, which is the correct model, and $\theta_{b}\neq q$, with $\theta_{b}<1$. The log-likelihood is given by
@@
Now suppose that in every step $t$ the Bayesian learner performs $\eta$-noisy updates where it over- or under-weights the data compared to the prior with a small excess weight $\eta_{t}$, where $\eta_{t}$ has zero mean. Specifically, for a parameter $\eta>0$, $\eta_{t}=\eta$ with probability $1/2$ and $\eta_{t}=-\eta$ otherwise. We find that even a tiny error with zero mean can have a significant impact, essentially breaking the learning process.
@@
For any $\eta>0$, the Bayesian learner with $\eta$-noisy updates does not converge.

### v2

@@ 5 The Regret of Bayesian Learners
In this section we further analyze the relationship between an agent’s regret and long-term survival in market dynamics, and use this analysis to derive the regret level of Bayesian learners. This relationship depends on the competition. In particular, we saw in Theorem 3.1 that the learners who survive (have a positive expected wealth level in the limit) are only those who maintain a constant regret gap relative to the best competitor in the market.
@@
Note that this agent uses information regarding future realizations of random states, which is not available to real agents in a stochastic environment. Next, we derive the regret level of using the correct distribution of states $q$ as the investment strategy. The proof is given in Appendix A; it uses our analysis from Section 3 and a result from information theory on the relative entropy between the underlying distribution $q$ and the empirical realization of states.
@@
We can now also state the following corollary.
@@
An agent indexed by $q$ using the correct distribution $q$ survives when competing against a magic agent. This is since by Theorem 5.2, the strategy $q$ yields constant expected regret $R$, and so the regret difference compared to $\hat{q}$—which by definition yields zero regret—is constant as well. By Theorem 3.1, this implies that agent $q$ survives.
Our analyses above and in the preceding sections lead to the following characterization of the regret of Bayesian learners and the survival conditions for other learning agents competing with them in the market. The proof is given in Appendix A.
@@
The next lemma provides an expression for the regret of a player as a function of the entropy relative to strategy $q$, using our notation $R$ from Theorem 5.2 for the expected regret of strategy $q$. This expression will be useful in the following Section on imperfect Bayesians.
@@
The expected regret of a constant strategy sequence (that is, $\alpha^{n}_{1:T}$ such that $\alpha^{n}_{t}=\alpha^{n}$ for all $t$) is given by $\operatorname{\mathbb{E}}[R^{T}(\alpha^{n}_{1:T})]=T\cdot I_{q}(\alpha^{n})+R.$
@@ 6 Imperfect Bayesians
In the preceding sections, we saw that a perfect Bayesian is optimal in the sense that it survives almost surely in the market and drives out any player with regret increasing over time. In this section, we explore a scenario where a Bayesian learner suffers small errors, either in its set of prior distributions or in its update rule. We find that, while perfect Bayesians are optimal, Bayesian learning is also very fragile; for example, it is key that the correct probability distribution is one of the models they consider. Specifically, we show that an imperfect Bayesian would eventually vanish from the market, either against the state distribution $q$ or in competition with any player with sub-linear regret. This holds even if the Bayesian’s errors are very small and have zero mean.
@@ 6.1 Bayesian Learning with Inaccurate Priors
Note that this contrasts with the case of Theorem 5.4, where a perfect Bayesian dominates the market in competition with any learners with regret increasing in time. The proof is in Appendix A.
@@
Our goal now is to analyze the survival time during which the inaccurate Bayesian may still maintain a significant market share, and to understand how this survival time is related to the level of error $\epsilon$. This, of course, also depends on the best competitor and how fast they converge.
To estimate this relation and provide an upper bound on the survival time of an inaccurate Bayesian learner, we consider a player using $q^{\prime}$ with $TV(q^{\prime},q)=\epsilon$ (having already converged to this distribution), competing against a second player playing a dynamic strategy with expected regret level increasing as $f(t)$ that is sub-linear in $t$.
First, we would like to express the first player’s error in terms of entropy (i.e., KL-divergence with $q$). We can bound the entropy using Pinsker’s inequality [78] for a lower bound and its inverse for an upper bound, were the latter holds for finite distributions with full support, as in our case. We have (recall that $\Delta=\min_{s}q_{s}$),
@@
To maximize the survival time, we pick the distribution $q^{\prime}$ to be the one with the smallest regret, which is when $I_{q}(q^{\prime})$ is the smallest possible: $I_{q}(q^{\prime})=2\epsilon^{2}$. So we have [^7] If the inaccurate agent uses the worst $q^{\prime}$ with error $\epsilon$, the regret calculation is similar, but with $\epsilon$ rescaled by a factor of $1/\sqrt{\Delta}$, using the inverse of Pinsker’s inequality: $2\epsilon^{2}\tau/\Delta+R$. The smallest-regret $q^{\prime}$ is used for the bound. $\mathbb{E}[R^{T}(q^{\prime}_{1:T})]=2\epsilon^{2}T+R$. By Equation (16), the regret difference between two players is equal to the log of their wealth ratios plus a constant. By comparing regret levels between the agents, we obtain a bound on the typical survival time $\tau$ up to which the inaccurate player holds, in expectation, more than half the market value; beyond this point, their expected share is less and continues decaying to zero: [^8] An alternative question one may ask is: how accurate does my prior need to be to allow retaining a share of the market for $T$ steps against the competitor? For this, one can invert the equation and get: $\epsilon=\sqrt{(f(T)-R)/T}$.
@@
Remark: Note that $f(t)$ is the competitor’s actual regret. Regret bounds for known learning algorithms (e.g., bandit algorithms such as UCB (see, e.g., [85]) or gradient-descent and second-order methods like those described in [61]) are typically worst-case bounds. Estimating the expected regret (or the most likely one) in game dynamics in general, or specifically in our investment scenario, is an interesting open problem.
For the special cases that the competitor has constant regret (for example, an accurate Bayesian), or the competitor has a $\log T$ or $\sqrt{T}$ regret level, we get the following results.
@@
For a competitor with logarithmic regret, we get an equation of the form $\tau=c_{1}\cdot\exp(c_{2}\epsilon^{2}\tau)$. For small errors $\epsilon$, by expanding the exponent, we have the following bound:
@@
Finally, in competition against a learner with regret $f(T)=\sqrt{T}$ the survival time is longer:
@@
The interpretation of this analysis naturally depends on the time scale of interest. When $T$ is large, which is our main focus (e.g., when trade occurs at high frequency or if investments are long-term), only long-term survival matters. In this case, a Bayesian with an inaccurate prior will eventually vanish in competition with any learner who converges to the truth (e.g., a regret minimizer). However, when transient dynamics are also relevant, it is possible that a Bayesian with a prior that is inaccurate but relatively close to the true distribution may retain a significant market share for some time before ultimately vanishing. See Section 8 for an example.
@@ 6.2 Bayesian Learning with Noisy Updates
Next, we consider a different type of imperfect Bayesian learner that does have $q$ in its prior, but in every step performs slightly inaccurate “trembling hand” updates. Also here, we demonstrate that Bayesian learning is fragile, even when the correct distribution lies in the support. To model this, we define a noisy Bayesian learner as one who at each step, either slightly over-weights the current observation or slightly over-weights its current prior, such that, in expectation, both the data and the prior receive the correct weights in every step (i.e., the errors in weight have zero mean).
For concreteness, consider the following scenario of a learner attempting to learn a distribution of states. Suppose that there are two states, $s_{t}=0$ with a fixed probability $q\in(0,1)$, and $s_{t}=1$ otherwise. The learner considers two models: $\theta_{a}=q$, which is the correct model, and $\theta_{b}\neq q$, with $\theta_{b}<1$. The log-likelihood is given by
@@
Now suppose that in every step $t$ the Bayesian learner performs “$\eta$-noisy updates” where it over- or under-weights the data compared to the prior with a small excess weight $\eta_{t}$, where $\eta_{t}$ has 00 mean. Specifically, for a parameter $\eta>0$, $\eta_{t}=\eta$ with probability $1/2$ and $\eta_{t}=-\eta$ otherwise. We find that even a tiny zero-mean error has a significant impact, essentially breaking the learning process.
@@
For any $\eta>0$, the Bayesian learner with $\eta$-noisy updates does not converge, and therefore incurs regret linear in $T$.
The idea of the proof is to show that, with high probability, the learner has a systematic drift toward over-weighing new observations, despite the symmetry of the errors around zero. As a result, the learner’s beliefs fail to converge and continue to fluctuate in response to recent random events, leading to linear regret. The full proof appears in Appendix A.
7 Robust Bayesian Updates
In the above analyses, we saw that the fast convergence of Bayesian learners is both a strength and a weakness. On the one hand, when a Bayesian converges to a correct model, it achieves constant regret and survives even against a fully informed competitor. On the other hand, if it converges to an incorrect model, it suffers linear regret (relative to the best fixed strategy in hindsight) and vanishes from the market. This strong convergence becomes especially problematic when the data-generating process undergoes distribution shifts.
In this section, we propose a simple method that naturally retains the fast convergence of Bayesian learning at early stages while regularizing its update to ensure robustness under distribution shifts. The method provides provable regret guarantees and improves adaptability. We consider a setting in which the data-generating distribution may shift over time. We first show that, while standard Bayesian learners may eventually adapt, their regret against the true sequence of distributions after a distribution shift can be linear in the total time $T$.
Proposition 7.1.
Consider a market in which the data-generating distribution $q$ may shift over the course of $T$ steps among a finite set $Q$ of distributions. A Bayesian learner with a finite prior that assigns positive probability to each model in $Q$ may incur regret linear in $T$ compared to a benchmark that knows the correct sequence of distributions in advance, even if there is only a single distribution shift during the $T$ steps.
The proof is in Appendix A. The Bayesian’s fast convergence is a feature if the Bayesian puts positive prior probability on a fixed, correct model of the data-generating process and for asymptotic results this all that matters; the prior is otherwise irrelevant. But if the data generating process shifts over time the Bayesian’s posterior at the (unknown) time of a shift becomes the prior. So intermediate-term properties matter. In the intermediate term, a Bayesian who has nearly converged to a past model learns slowly, as the log odds of the new model to the past one are extreme and the expected shift in these log odds is the fixed relative entropy between the two models. With repeated shifts, the Bayesian thus spends a large amount of time near models that are incorrect and accumulates regret that can be linear.
In contrast, there are no-regret learning algorithms that are known to handle distribution shifts more gracefully [60, 70]. In particular, for the portfolio selection problem we consider, such algorithms can achieve regret that is logarithmic in $T$ and linear in the number $n$ of distribution shifts (see also [61], Sections 10.3 and 10.4). However, as we have seen, logarithmic regret is not sufficient for survival when competing against agents who invest according to the true distribution, or even against Bayesian learners.
The main observation of this section is that, while the exponential convergence that Bayesians achieve is sufficient to obtain constant regret—and thus ensures long-run survival against fully informed competitors—it is not necessary. We propose a simple and natural method that combines advantages of both approaches, making a step towards a best-of-both-worlds learning strategy. It guarantees constant regret in stationary settings, like standard Bayesian learning, thus outperforming no-regret learners in such cases. At the same time, it guarantees logarithmic regret under distribution shifts, surviving against no-regret learners in settings where adaptability is essential.
Technically, our method stays close to Bayesian learning. We use a regularized version of the Bayesian update rule that initially behaves like standard Bayesian learning, but slows convergence just enough to keep alternative hypotheses alive. This enables fast recovery after shifts.
Robust Bayesian Update: The Robust Bayesian Update process we propose uses a simple and efficient modification of standard Bayesian learning. Initially, the learner sets uniform prior weights $\lambda^{1}_{0},\dots,\lambda^{K}_{0}$. At each time $t\geq 0$, the learner uses the current weights $\lambda_{t}$ as its investment strategy. Then, after observing the outcome, it updates beliefs with an added regularization term $\epsilon_{t}=t^{-2}$:
$\lambda_{t}\rightarrow\text{ (Bayes update) }\rightarrow\tilde{\lambda}_{t+1}\rightarrow\lambda_{t+1}=\frac{\tilde{\lambda}_{t+1}+\epsilon_{t}}{1+K\epsilon_{t}}.$ (26)
where $K$ is the number of hypotheses. That is, using the observed outcome at time $t$, the learner applies the standard Bayesian update to obtain tentative weights $\tilde{\lambda}_{t+1}$. It then adds regularization and normalizes to get weights $\lambda_{t+1}$ in the simplex. This simple process satisfies the following guarantee:
Theorem 7.2.
Let $Q$ be the set of models in the learner’s support. Suppose an adversary selects a sequence of intervals ($\big{(}\{T_{i},q_{i}\}\big{)}_{i=1}^{n}$, where $q_{i}\in Q$ and $T_{i}$ is the duration of the $i$-th interval during which $q_{i}$ is the data-generating distribution. Let $T=\sum_{i=1}^{n}T_{i}$ denote the total time. Then the regret of the Robust Bayesian Update, relative to a benchmark of fixed investment strategies in each interval, knowing the intervals in advance, is $O(n\log T)$ with high probability. In particular, when $n$ is constant, the regret is logarithmic in $T$.
Moreover, if there are no distribution shifts (i.e., when $n=1$), the regret remains bounded by a constant at all times with high probability.
The intuition behind the proof is that when the weight of an incorrect model is large, it decays exponentially fast, as in the case of a standard Bayesian learner. When it is small—approaching the scale of the regularization term—it no longer contributes significantly to regret. Thus, the regret under a stationary process remains bounded. After a distribution shift, all model weights are of at least the size of the regularization level, which prevents suppressing the alternatives at an exponential rate and keeps them “alive.” This is then followed by exponential convergence, leading to regret that grows only logarithmically in $T$. The proof of the theorem is in Appendix A.
8 Simulations
(a) Wealth Dynamics. In green: player 2’s wealth at each step. Solid lines show sliding window averages over the recent 10,000 steps for each $t$.
(b) Wealth distribution in the second half of the $10^{6}$-step simulation. The distribution shown in blue is the wealth of player 1 and in green is of player 2.
(c) Wealth distribution in the 10,000 initial steps of the dynamic. The distribution shown in blue is the wealth of player 1 and in green is of player 2.
Figure 1: Example of wealth dynamics in competition between an inaccurate Bayesian with an error in its support (Agent 1) learner and no-regret learner (Agent 2) who converges to the correct state distribution $q$.
Imperfect Bayesian Learning: Consider the following simple example. There is market with two assets where agent 1 is a Bayesian learner and agent 2 is a UCB learner, starting from equal wealth levels. [^9] We use UCB for simplicity. The main factor here is the different regret rates and not the particular algorithm. The distribution of states is $q=(0.7,0.3)$, i.e., state 1 occurs with probability $0.7$, and state 2 otherwise. The Bayesian learner has a prior consisting of three hypotheses: $(0.8,0.2)$, $(0.9,0.1)$, $(0.3,0.7)$, where the first hypothesis is closest to $q$ in relative entropy, but still has an error. The UCB learner considers a similar set of distributions as its action set: $(0.7,0.3)$, $(0.9,0.1)$, $(0.3,0.7)$, but the first of these being the correct distribution $q$. The example demonstrates the trade-off between short-term gains for the inaccurate Bayesian from converging quickly to a distribution close to the correct one, and the long-term dominance of the learner who converges to the correct distribution.
(a) Two Bayesian learners.
(b) Two no-regret learners.
(c) Bayesian vs. no-regret learner.
Dynamics with a Constant Regret Difference: Next, we consider an example of two Bayesian learners, both with priors including the correct distribution of states. By Theorem 5.4, their regret levels can differ only by an additive constant, as both are perfect Bayesians with constant regret. Consequently, these agents are bound to coexist in the market. The example demonstrates that while the quality of the priors—provided they include a correct model—does not affect survival, the initial learning phase, which depends on the prior, does influence the steady-state partition of wealth between the agents. In this example, as in the previous one, there are two states, with the underlying distribution given by $q=(0.7,0.3)$. Agent 1 has a prior that includes the correct distribution and two other models: $(0.8,0.2),(0.7,0.3),(0.6,0.4)$. Agent 2’s prior includes one additional model: $(0.8,0.2),(0.7,0.3),(0.6,0.4),(0.5,0.5)$. Both agents start with a uniform prior over the options they consider. Thus, agent 2 places smaller weight initially on the correct model than does agent 1. Figure 2(a) depicts the wealth dynamics of these learners, starting from equal wealth levels.
9 Conclusion
As markets become increasingly populated by learning agents, understanding which learning approaches succeed in competitive environments becomes a foundational question. Our work brings together perspectives from economics and computer science by analyzing the survival and performance of Bayesian and no-regret learners in complete asset markets. We connect the notions of regret and survival in a competitive market, and show that regret minimization, while robust and widely applicable, does not guarantee survival: a Bayesian investor with correct support and proper updates can drive out no-regret learners, even if their regret is vanishing. Conversely, Bayesian learning is fragile—slight misspecification in priors or slight mistakes in belief updates may lead to linear regret and eventual extinction from the market.
These insights motivate the study of more robust ways to utilize Bayesian learning and combine advantages from both approaches. We proposed and analyzed a simple regularized variant of Bayesian updating that preserves fast convergence in stationary environments while ensuring adaptability to distribution shifts. The resulting method achieves constant regret in stable settings and only logarithmic regret when the environment changes, thereby dominating existing methods and offering a simple, practical step toward a best-of-both-worlds approach. Our findings highlight the speed of learning as a crucial determinant of survival and show that even constant factors in regret rates matter in competitive markets. We view the development of robust Bayesian learning methods, guided by principles from online learning theory, as a promising direction for future work.
Another interesting direction for future work is to consider the impact of competing learners on aggregate outcomes such as equilibrium prices. Both long run security prices, determined by who survives, and intermediate term security prices, determined by the evolution of investment rules and wealth shares, could be analyzed. Our focus has been on the survival of individual agents and they determine long run prices. But intermediate term prices also depend on how agents respond to the feedback they receive over time from markets and how wealth shares evolve between these agents. Both Bayesian and standard no-regret learning cause agents to invest more in assets that pay off and wealth share increases for agents who invest relatively more in these assets. So intermediate term security prices respond positively to past payoffs. These learning and wealth share evolution effects induce positive serial correlation—or momentum—in security prices.
Acknowledgments
We thank the anonymous referees at EC 2025 for useful feedback on an earlier version of the paper. Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068, and ONR MURI grant N000142412742.
APPENDIX
Appendix A Proofs
We provide in the following the proofs for the theorems stated in the main text.